OpenStudy (anonymous):

Solve the equation. 7x^4-33x^2+26=0 Please!! :)

OpenStudy (anonymous):

let x^2 = t now 7t^2 - 33 t + 26 = 0 one root you get as 1 other = 26/7 which are values for t = x^2 now x ^2 = 1 and 26 /7 x = +1,-1, \[+\sqrt{26/7} ,, and -\sqrt{26/7}\] which are answers to this equation